please help? Thank you so much :)

...He said the banking system could lose deposits in several ways. One channel in recent quarters has been domestic fund mAssets Bank B Liabilities and equity ESF AUD 100m Deposits AUD 700m Loans AUD 900m Deposits Mr Red. AUD 100m Equity AUD 200m
Answer #1

Part A

                                         OZ Bank

ASSETS                                                 LIABILTIES & EQUITY

ESF                         AUD 150M                    Deposit                         AUD 810M

Loans                       AUD 850M                   Household                     AUD 50M

                                                                       Equity                           AUD 40M

                                       

                                      FUND MANAGER

ASSETS                                                               LIABILTIES & EQUITY

Bonds                      AUD 100M                         Equity                    AUD 100M

                                     

                                              BANK B

ASSETS                                                                 LIABILTIES & EQUITY   

ESF                        AUD 100M                           Deposits               AUD 700M

Loans AUD 900M    Deposit Mr.RED AUD 100M

Equity                 AUD 200M                

Loans to deposit ratio for bank B remains unchanged. for bank Oz

Earlier =Loans/ Deposit = (850/910)*100= 93.41%

Now =Loans/ Deposit =(850/810)*100= 104.94%

PART B

                                            BANK B

ASSETS                                                                 LIABILTIES & EQUITY   

ESF                        AUD 100M                           Deposits               AUD 700M

Loans AUD 800M                 Equity                 AUD 200M       

Loans to deposit ratio changes.

earlier =Loans/ Deposit =(900/800)*100= 112.50%

Now =Loans/ Deposit =(800/700)*100= 114.29%

PART C

                                         OZ Bank

ASSETS                                                 LIABILTIES & EQUITY

ESF                         AUD 150M                    Deposit                         AUD 860M

Loans                       AUD 850M                   Household                     AUD 50M

BONDS                   AUD 50M                    Equity                           AUD 40M